40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 7912     Entry time: Thu Jan 17 11:01:19 2013
Author: Jenne 
Type: Update 
Category: Alignment 
Subject: Yesterday's alignment work 

[Jamie, Jenne, Manasa]

Yesterday's goal was to get the input beam centered on the PRM, the BS and ETMY simultaneously. 

Steve helped us remove the ETMY door first thing in the morning.  We then iterated with TT1, MMT1 and TT2 to try to get the beam centered on all the optics.  We were using MMT1 instead of TT1 for a while, so that we could keep TT1 in the center of its range, so that we had more range to use once we pump down.  Also, at one point, the beam was high on PRM, centered on BS, and high on ETMY, so Jamie poked PR3 a little bit.  This helped, although we closed up for lunch / group meeting soon after, so we didn't finalize any alignment stuff.

We decided to leave the rest of the full IFO alignment alone until after the PRM-flat test.

ELOG V3.1.3-